# Exploring Large-Volume GAGG Scintillators for use in MeV Gamma-Ray Astrophysics

### v2

@@ 3 High-Energy Gamma-Ray Response of the Calorimeter
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peak, likely due to the low full-energy efficiency and full-containment of the ensuing spallation. The experimental response matches the simulated one, presented in A.
@@ 5 Conclusion and Future Work
We are currently developing a larger $8\times 8$ array calorimeter using the $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals. In addition, we are developing SiPM readouts that field two different SiPM species such that their different gains will the dynamic energy range of the readout [26]. The ‘dual-gain’ SiPMs are currently under fabrication by Fondazione Bruno Kessler [27].

### v3

@@ 3 High-Energy Gamma-Ray Response of the Calorimeter
Fig. 8 plots all the acquired spectra from the HIGS campaign. Full energy peaks are clearly visible for the 2, 5, and 7 $\mathrm{MeV}$ beams. Also featured are the single and double escape peaks. However, runs with 15 and 25 $\mathrm{MeV}$ do not present any evidence of escape peaks, likely due to the low full-energy efficiency and full-containment of the ensuing spallation. The experimental response matches the simulated one, presented in A.
@@ 5 Conclusion and Future Work
We are currently developing a larger $8\times 8$ array calorimeter using the $1\times 1\times 8\ \mathrm{cm}^{3}$ crystals. In addition, we are developing SiPM readouts that field two different SiPM species such that their different gains will increase the dynamic energy range of the readout [26]. The ‘dual-gain’ SiPMs are currently under fabrication by Fondazione Bruno Kessler [27].